# Playing with Voices: Tabletop Role-Playing Game Recordings as a Diarization Challenge

### v1

@@ 2.3 Reference datasets: AMI and ICSI corpora
As one of the applied diarizers, pyannote.audio Bredin (2023) (see section 4.2) was trained on the AMI corpus, it should perform particularly well on it. This means a comparison with the AMI corpus keeps our findings of pyannote.audio conservative. Therefore, we only used the test dataset which pyannote.audio was not trained on (16 files) from the AMI corpus on pyannote.audio. These test files contain and of audio.
@@ 4 Data and methods
This section provides an overview about how we created our TTRPG dataset and how we applied a diarizer on the audio files. Our TTRPG dataset was compiled using publicly available videos and subtitles. As this paper is a proof of concept and the data are not shared to any third-party, we had no ethical concerns in experimenting with the audio files (see section 8). All scripts for the data processing and analyses, and links to the source videos we compiled the TTRPG data from are publicly available.111https://github.com/LiRem101/playing-with-voices The processed annotated transcripts are available from the authors upon request.
@@ 4.1 Creating TTRPG dataset
Five of the six campaigns consisted of battle and role-playing scenes (see section 2), one contained roleplay scenes only. We selected one hour of roleplay and, if available, one hour of battle audio from each campaign. We identified the onset of the roleplay portion using the first interaction between two player characters, and the onset of the battle portion when the GM signals the start of the battle. One campaign was removed, as cursory inspection revealed high subtitle inaccuracies. This resulted in 21 files with and of battle, and and of roleplay. The speakers speak American English. We could determine the age of of the speakers (mean: years, min: , max: ). This was determined by checking the people’s Wikipedia entries or their public social media channels and using the day the respective video was uploaded as the reference date.
The subtitles were used to create the ground truth diarization files. Some included the information about who spoke what, for others this was added manually. We applied forced alignment to the audio files and subtitles using Wav2Vec2FABundle Baevski et al. (2020). Combined with the knowledge of who said which words, we created diarization ground truth files by looking at the time stamps of each word and treating a gap of or less between two words of the same speaker as one utterance. While how short pauses are handled is not described for AMI and ICSI, a different approach from ours for the reference datasets should not influence our results, because we ignore around the start and the end of an utterance in the evaluation (see section 5.2).
A shortcoming of creating diarization ground truth this way is that overlapping speech as well as utterances that are not included in the subtitles files are not taken into account. These are further investigated in section 5.3. To overcome this, we manually annotated ten minutes of one role-playing audio manually without help of subtitle files or forced alignment. This annotation contains the information of whether the speech was in-character or not. As this is a proof of concept, the annotations was created by only one person (one of the authors, ANON) with the help of spectrograms, therefore inter-annotation-agreement was not evaluated.
@@ 5 Results
In this section, we present our findings about the diarization of TTRPG audio files compared to the AMI and ICSI dataset. We used pyannote.audio (v. 3.1.0) and wespeaker (v. 1.2.0) to diarize 21 TTRPG audio files, the 171 (16 for pyannote.audio) AMI audio files, and the 75 ICSI audio files. The evaluation was done using pyannote.metrics (v. 3.2.1) Bredin (2017). We applied Mann-Whitney U tests Mann and Whitney (1947) to some result datasets to check whether the distributions of the datasets were significantly different, using a significance threshold of . In this section, we only comment if the differences were significant or not. The exact values of the Mann-Whitney U tests can be found in appendix A.
@@ 5.1 Amount of detected speakers
The hypothesis of the diarizer finding more speakers in the TTRPG dataset was not supported. We measured the relative amount of detected speakers divided by the amount of actual speakers. The results of pyannote.audio can be seen in fig. 1. It found on average for the AMI dataset, for the ICSI dataset, and for the TTRPG dataset. This difference was not statistically significant in a two-sided Mann-Whitney U test (table 2, appendix A). The results of wespeaker can be seen in fig. 2. It found on average for the AMI dataset, for the ICSI dataset, and for the TTRPG dataset. A two-sided Mann-Whitney U test Mann and Whitney (1947) showed significant differences between all distributions (table 2, appendix A).
Considering AMI and ICSI, the relation of detected speakers against the actual number of speakers is significantly smaller for wespeaker than for pyannote.audio (table 2, appendix A), and wespeaker’s predictions are closer to the actual value. However, wespeaker drastically underestimates the number of speakers in TTRPG.
@@ 5.2 Diarization errors
All average error rates and Mann-Whitney U test results used in this section to check for significant differences can be found in table 3 (appendix A).
Both pyannote.audio and wespeaker had the highest average DER while diarizing the TTRPG, with (pyannote.audio) and (wespeaker) respectively (figs. 6 and 7, appendix B). The results on the other datasets were significantly lower. pyannote.audio reached an average of on AMI and on ICSI, while wespeaker had an average of on AMI and on ICSI (table 3, appendix A).
@@
The missed detection rates between TTRPG and AMI are not significantly different for pyannote.audio or wespeaker, while ICSI differs significantly from AMI and TTRPG for both diarizers (table 3, appendix A). The respective boxplots can be found in figs. 8 and 9, appendix B.
The TTRPG false alarm rates are significantly higher than the AMI rates for both pyannote.audio and wespeaker. The TTRPG dataset’s false alarm also differs significantly from the ICSI dataset, but is not significantly higher, since the ICSI dataset shows the highest false alarm rate of all datasets (table 3, appendix A). The distributions can be seen in figs. 10 and 11, appendix B.
As with both diarizers ICSI has higher false alarm but lower missed detection than AMI, we suspect that the differences between those two datasets root in a more careful annotation on ICSI’s reference files. This is backed up by fewer interjections in the ICSI reference files than in the AMI reference files (see section 5.3) and the fact that both diarizers showed this result.
However, the diarizers do not score well on missed detection or false alarm on the TTRPG dataset. We suspected that this was not due to the poor performance of the diarizer, but by the quality of the ground truth files. This is further evaluated in sections 5.3 and 5.4.
@@
Figures 3 and 4 show the confusion of the diarizers for the TTRPG dataset compared to the AMI and ICSI dataset. For the confusion, the TTRPG rates are significantly higher than the AMI rates and the ICSI rates for both diarizers, while the confusion of AMI and ICSI datasets does not differ significantly on pyannote.audio, but does differ on wespeaker (table 3, appendix).
To establish whether these performance differences are independent of the diarizer’s ability to detect the number of speakers (section 5.1), we repeated our diarization experiments with the actual number of speakers given to the system. This modification did not influence the error rates of missed detection and false alarm. While we expected that it would decrease the confusion error, it in fact led to an increase of the AMI and ICSI dataset confusion on pyannote.audio, to a point where the difference between both datasets was no longer significant (AMI: , , ICSI: , , TTRPG: , ). On wespeaker, the average confusion rates changed for AMI and ICSI if we gave the number of speakers to the diarizer (AMI: , , ICSI: , , TTRPG: , ). However, the AMI and ICSI confusion stayed significantly smaller than the TTRPG confusion (table 3, appendix A). The respective boxplots can be found at figs. 12 and 13, appendix B.
@@ 5.3 Weaknesses of the TTRPG reference files
The TRPG reference files were created with subtitles and forced alignment, as described in section 4.1. While with this approach the ground truth can be efficiently obtained, it does negatively influence its quality because it depends on the quality of the subtitles and of the forced alignment process. This has been especially evident when we had to exclude one of the chosen campaigns, after noticing its gross error rates and finding that its subtitle was far from being verbatim (see section 4.1).
Even after excluding campaigns with gross transcription issues, the approach of using subtitles had two problems. First, subtitles often lacked filler words like “uhmm” or “ehh” and paraphrased the meaning of words. Secondly, the use of subtitles and forced alignment ignored overlapping speech by multiple speakers. It either ignores all but one speaker or appears as if the words would have been said after each other. Both problems would lead to missing speech in the ground truth, which can result in an apparently higher false alarm rate we observed in section 5.2 compared to AMI. To verify this further, we evaluated the properties of the TTRPG dataset and compared it to AMI and ICSI.
@@ 5.4 Manually annotated reference file
The resulting error rates can be seen in table 1, compared to the error rates from the automatically created reference files. The manually created reference files result in overall smaller error rates when pyannote.audio was used. While the confusion rate only decreased by , the false alarm decreased by and the missed detection by . If the number of speakers was given to the diarizer upfront, the rates decreased by , , and respectively. For wespeaker, all error rates but missed detection got smaller. The confusion rate decreased by ( if speaker number was given) and the false alarm by (). The missed detection rate raised by ().
@@ 5.5 Detailed error analyses
To further understand the errors of the diarizer, we examined the confusion matrices for the manually created reference (see fig. 5; the rows of the matrices are normalized). For the matrices, we only examined regions that have been labeled as speech in the reference and the prediction, ignoring false alarm and missed detection. The manually created reference contained information whether speech was “in-character” or “descriptive”. We differentiated in the matrices between these two ways of speaking for every person.
Figure 5(a) (left) is the confusion matrix when pyannote.audio has not been given the number of actual speakers. Two male speakers (M1 and M2) have been “merged” and are mostly mapped to the predicted SP_01. In this audio file, M0 has had the position of the GM, meaning that the speaker had not one in-character voice but several (see section 2.2). This seems to lead to wrong predictions by the diarizer, such as that in-character of M0 “creates” a new speaker SP_04. Additionally, SP_0 has been assigned to many other actual speakers, especially M1 and F0.
@@
Figure 5(a) (right) is the confusion matrix when pyannote.audio has been given the number of actual speakers. All male speakers are mapped to the same predicted speaker. Two extra speakers (SP_04 and SP_03) consist of some of M1’s utterances and the in-character voice of F0. Again, the change of voice leads to a higher scattering of detected speakers. The average entropy over all rows, descriptive voices and in-character voices are , , and respectively.
Figure 5(b) shows the confusion matrix when wespeaker has been given the number of actual speakers. A confusion matrix without the speaker number given has not been created, since wespeaker predicted only one speaker in this case. This aligns with our findings that wespeaker finds small numbers of speakers for the TTRPG files (see section 5.1). The average entropy over all rows, descriptive voices and in-character voices are , , and . In the case of wespeaker, the in-character voices are not more scattered than the descriptive voices. However, the scattering is higher than pyannote.audio’s overall.
@@ 6 Conclusion
This paper aimed to provide a proof of concept that TTRPG audio files serve as a complex but natural challenge to diarization systems. We were able to show that pyannote.audio’s and wespeaker’s speaker confusion increased for TTRPG compared to AMI and ICSI audio files (see sections 5.2 and 5.4), which we consider to be similar except for the unique TTRPG properties (see section 2.3). We found evidence that a low-resource method of annotating a TTRPG dataset does not conceal the fact that a diarizer gets confused by TTRPGs’ properties. Additionally, we found that it could be advantageous to annotate whether utterances are in-character or descriptive, to be able to evaluate the diarization performance in more depth. pyannote.audio’s and wespeaker’s confusion increased for TTRPG audios. Nevertheless, the relative amount of speakers pyannote.audio found compared to the amount of speakers that actually are in the audio file did not change. In case of wespeaker we found that the number decreased, indicating that its clustering algorithm gets confused by people changing their voices frequently (see section 5.1).
@@ 7 Limitations
This work presents a proof of concept which needs further investigation to ensure our findings can be generalized. Other SOTA diarizers should be tested on top of pyannote.audio and wespeaker. Additionally, we only compared the diarization performance to the AMI and ICSI corpora and had a relatively small sample size. As mentioned in section 5.3, the ground truth for the TTRPGs is imperfect, however as shown in section 5.4 we showed evidence that the imperfect ground truth do not compromise our core findings. Our claim concerning the reduction in confusion error rates for manually aligned audio in Section 5.4 needs to be further validated statistically. This can be done in a larger study that examines multiple sets of manual and automatic annotated audio.

### v2

@@ 2.3 Reference datasets: AMI and ICSI corpora
As one of the applied diarizers, pyannote.audio Bredin (2023) (see section˜4.2) was trained on the AMI corpus, it should perform particularly well on it. This means a comparison with the AMI corpus keeps our findings of pyannote.audio conservative. Therefore, we only used the test dataset which pyannote.audio was not trained on (16 files) from the AMI corpus on pyannote.audio. These test files contain and of audio.
@@ 4 Data and methods
This section provides an overview about how we created our TTRPG dataset and how we applied a diarizer on the audio files. Our TTRPG dataset was compiled using publicly available videos and subtitles. As this paper is a proof of concept and the data are not shared to any third-party, we had no ethical concerns in experimenting with the audio files (see section˜8). All scripts for the data processing and analyses, and links to the source videos we compiled the TTRPG data from are publicly available.111https://github.com/LiRem101/playing-with-voices The processed annotated transcripts are available from the authors upon request.
@@ 4.1 Creating TTRPG dataset
Five of the six campaigns consisted of battle and role-playing scenes (see section˜2), one contained roleplay scenes only. We selected one hour of roleplay and, if available, one hour of battle audio from each campaign. We identified the onset of the roleplay portion using the first interaction between two player characters, and the onset of the battle portion when the GM signals the start of the battle. One campaign was removed, as cursory inspection revealed high subtitle inaccuracies. This resulted in 21 files with and of battle, and and of roleplay. The speakers speak American English. We could determine the age of of the speakers (mean: years, min: , max: ). This was determined by checking the people’s Wikipedia entries or their public social media channels and using the day the respective video was uploaded as the reference date.
The subtitles were used to create the ground truth diarization files. Some included the information about who spoke what, for others this was added manually. We applied forced alignment to the audio files and subtitles using Wav2Vec2FABundle Baevski et al. (2020). Combined with the knowledge of who said which words, we created diarization ground truth files by looking at the time stamps of each word and treating a gap of or less between two words of the same speaker as one utterance. While how short pauses are handled is not described for AMI and ICSI, a different approach from ours for the reference datasets should not influence our results, because we ignore around the start and the end of an utterance in the evaluation (see section˜5.2).
A shortcoming of creating diarization ground truth this way is that overlapping speech as well as utterances that are not included in the subtitles files are not taken into account. These are further investigated in section˜5.3. To overcome this, we manually annotated ten minutes of one role-playing audio manually without help of subtitle files or forced alignment. This annotation contains the information of whether the speech was in-character or not. As this is a proof of concept, the annotations was created by only one person (one of the authors, ANON) with the help of spectrograms, therefore inter-annotation-agreement was not evaluated.
@@ 5 Results
In this section, we present our findings about the diarization of TTRPG audio files compared to the AMI and ICSI dataset. We used pyannote.audio (v. 3.1.0) and wespeaker (v. 1.2.0) to diarize 21 TTRPG audio files, the 171 (16 for pyannote.audio) AMI audio files, and the 75 ICSI audio files. The evaluation was done using pyannote.metrics (v. 3.2.1) Bredin (2017). We applied Mann-Whitney U tests Mann and Whitney (1947) to some result datasets to check whether the distributions of the datasets were significantly different, using a significance threshold of . In this section, we only comment if the differences were significant or not. The exact values of the Mann-Whitney U tests can be found in appendix˜A.
@@ 5.1 Amount of detected speakers
The hypothesis of the diarizer finding more speakers in the TTRPG dataset was not supported. We measured the relative amount of detected speakers divided by the amount of actual speakers. The results of pyannote.audio can be seen in fig.˜1. It found on average for the AMI dataset, for the ICSI dataset, and for the TTRPG dataset. This difference was not statistically significant in a two-sided Mann-Whitney U test (table˜2, appendix˜A). The results of wespeaker can be seen in fig.˜2. It found on average for the AMI dataset, for the ICSI dataset, and for the TTRPG dataset. A two-sided Mann-Whitney U test Mann and Whitney (1947) showed significant differences between all distributions (table˜2, appendix˜A).
Considering AMI and ICSI, the relation of detected speakers against the actual number of speakers is significantly smaller for wespeaker than for pyannote.audio (table˜2, appendix˜A), and wespeaker’s predictions are closer to the actual value. However, wespeaker drastically underestimates the number of speakers in TTRPG.
@@ 5.2 Diarization errors
All average error rates and Mann-Whitney U test results used in this section to check for significant differences can be found in table˜3 (appendix˜A).
Both pyannote.audio and wespeaker had the highest average DER while diarizing the TTRPG, with (pyannote.audio) and (wespeaker) respectively (figs.˜6 and 7, appendix˜B). The results on the other datasets were significantly lower. pyannote.audio reached an average of on AMI and on ICSI, while wespeaker had an average of on AMI and on ICSI (table˜3, appendix˜A).
@@
The missed detection rates between TTRPG and AMI are not significantly different for pyannote.audio or wespeaker, while ICSI differs significantly from AMI and TTRPG for both diarizers (table˜3, appendix˜A). The respective boxplots can be found in figs.˜8 and 9, appendix˜B.
The TTRPG false alarm rates are significantly higher than the AMI rates for both pyannote.audio and wespeaker. The TTRPG dataset’s false alarm also differs significantly from the ICSI dataset, but is not significantly higher, since the ICSI dataset shows the highest false alarm rate of all datasets (table˜3, appendix˜A). The distributions can be seen in figs.˜10 and 11, appendix˜B.
As with both diarizers ICSI has higher false alarm but lower missed detection than AMI, we suspect that the differences between those two datasets root in a more careful annotation on ICSI’s reference files. This is backed up by fewer interjections in the ICSI reference files than in the AMI reference files (see section˜5.3) and the fact that both diarizers showed this result.
However, the diarizers do not score well on missed detection or false alarm on the TTRPG dataset. We suspected that this was not due to the poor performance of the diarizer, but by the quality of the ground truth files. This is further evaluated in sections˜5.3 and 5.4.
@@
Figures˜3 and 4 show the confusion of the diarizers for the TTRPG dataset compared to the AMI and ICSI dataset. For the confusion, the TTRPG rates are significantly higher than the AMI rates and the ICSI rates for both diarizers, while the confusion of AMI and ICSI datasets does not differ significantly on pyannote.audio, but does differ on wespeaker (table˜3, appendix).
To establish whether these performance differences are independent of the diarizer’s ability to detect the number of speakers (section˜5.1), we repeated our diarization experiments with the actual number of speakers given to the system. This modification did not influence the error rates of missed detection and false alarm. While we expected that it would decrease the confusion error, it in fact led to an increase of the AMI and ICSI dataset confusion on pyannote.audio, to a point where the difference between both datasets was no longer significant (AMI: , , ICSI: , , TTRPG: , ). On wespeaker, the average confusion rates changed for AMI and ICSI if we gave the number of speakers to the diarizer (AMI: , , ICSI: , , TTRPG: , ). However, the AMI and ICSI confusion stayed significantly smaller than the TTRPG confusion (table˜3, appendix˜A). The respective boxplots can be found at figs.˜12 and 13, appendix˜B.
@@ 5.3 Weaknesses of the TTRPG reference files
The TRPG reference files were created with subtitles and forced alignment, as described in section˜4.1. While with this approach the ground truth can be efficiently obtained, it does negatively influence its quality because it depends on the quality of the subtitles and of the forced alignment process. This has been especially evident when we had to exclude one of the chosen campaigns, after noticing its gross error rates and finding that its subtitle was far from being verbatim (see section˜4.1).
Even after excluding campaigns with gross transcription issues, the approach of using subtitles had two problems. First, subtitles often lacked filler words like “uhmm” or “ehh” and paraphrased the meaning of words. Secondly, the use of subtitles and forced alignment ignored overlapping speech by multiple speakers. It either ignores all but one speaker or appears as if the words would have been said after each other. Both problems would lead to missing speech in the ground truth, which can result in an apparently higher false alarm rate we observed in section˜5.2 compared to AMI. To verify this further, we evaluated the properties of the TTRPG dataset and compared it to AMI and ICSI.
@@ 5.4 Manually annotated reference file
The resulting error rates can be seen in table˜1, compared to the error rates from the automatically created reference files. The manually created reference files result in overall smaller error rates when pyannote.audio was used. While the confusion rate only decreased by , the false alarm decreased by and the missed detection by . If the number of speakers was given to the diarizer upfront, the rates decreased by , , and respectively. For wespeaker, all error rates but missed detection got smaller. The confusion rate decreased by ( if speaker number was given) and the false alarm by (). The missed detection rate raised by ().
@@ 5.5 Detailed error analyses
To further understand the errors of the diarizer, we examined the confusion matrices for the manually created reference (see fig.˜5; the rows of the matrices are normalized). For the matrices, we only examined regions that have been labeled as speech in the reference and the prediction, ignoring false alarm and missed detection. The manually created reference contained information whether speech was “in-character” or “descriptive”. We differentiated in the matrices between these two ways of speaking for every person.
Figure˜5(a) (left) is the confusion matrix when pyannote.audio has not been given the number of actual speakers. Two male speakers (M1 and M2) have been “merged” and are mostly mapped to the predicted SP_01. In this audio file, M0 has had the position of the GM, meaning that the speaker had not one in-character voice but several (see section˜2.2). This seems to lead to wrong predictions by the diarizer, such as that in-character of M0 “creates” a new speaker SP_04. Additionally, SP_0 has been assigned to many other actual speakers, especially M1 and F0.
@@
Figure˜5(a) (right) is the confusion matrix when pyannote.audio has been given the number of actual speakers. All male speakers are mapped to the same predicted speaker. Two extra speakers (SP_04 and SP_03) consist of some of M1’s utterances and the in-character voice of F0. Again, the change of voice leads to a higher scattering of detected speakers. The average entropy over all rows, descriptive voices and in-character voices are , , and respectively.
Figure˜5(b) shows the confusion matrix when wespeaker has been given the number of actual speakers. A confusion matrix without the speaker number given has not been created, since wespeaker predicted only one speaker in this case. This aligns with our findings that wespeaker finds small numbers of speakers for the TTRPG files (see section˜5.1). The average entropy over all rows, descriptive voices and in-character voices are , , and . In the case of wespeaker, the in-character voices are not more scattered than the descriptive voices. However, the scattering is higher than pyannote.audio’s overall.
@@ 6 Conclusion
This paper aimed to provide a proof of concept that TTRPG audio files serve as a complex but natural challenge to diarization systems. We were able to show that pyannote.audio’s and wespeaker’s speaker confusion increased for TTRPG compared to AMI and ICSI audio files (see sections˜5.2 and 5.4), which we consider to be similar except for the unique TTRPG properties (see section˜2.3). We found evidence that a low-resource method of annotating a TTRPG dataset does not conceal the fact that a diarizer gets confused by TTRPGs’ properties. Additionally, we found that it could be advantageous to annotate whether utterances are in-character or descriptive, to be able to evaluate the diarization performance in more depth. pyannote.audio’s and wespeaker’s confusion increased for TTRPG audios. Nevertheless, the relative amount of speakers pyannote.audio found compared to the amount of speakers that actually are in the audio file did not change. In case of wespeaker we found that the number decreased, indicating that its clustering algorithm gets confused by people changing their voices frequently (see section˜5.1).
@@ 7 Limitations
This work presents a proof of concept which needs further investigation to ensure our findings can be generalized. Other SOTA diarizers should be tested on top of pyannote.audio and wespeaker. Additionally, we only compared the diarization performance to the AMI and ICSI corpora and had a relatively small sample size. As mentioned in section˜5.3, the ground truth for the TTRPGs is imperfect, however as shown in section˜5.4 we showed evidence that the imperfect ground truth do not compromise our core findings. Our claim concerning the reduction in confusion error rates for manually aligned audio in Section 5.4 needs to be further validated statistically. This can be done in a larger study that examines multiple sets of manual and automatic annotated audio.